A mixture containing 1.12 L of H2 and 1.12 L of D2 at STP is taken insidea bulb connected to another bulb by a stopcock with a small opening. The second bulb is fullt evacuated, the stopcock is opened for a certain time and then closed. The first bulb is found to comtain 0.05g og H2 . Dtermine the percentage composition by weight of the gases in the second bulb.

Grade:12

1 Answers

Aman Bansal
12 years ago

Dear Abhijat,

Rate of diffusion ~ 1/ molecular mass
MD2 = 4
MH2= 2

RH2/RD2 = (2)^{1/2}

As temp and volume (of the gases in the second container) remain constant

R ~ Pressure of gases (P)

P ~ no. of moles (n) => R ~ n

nD2 = nH2/ (2)^1/2

Initial moles of H2 in 1st bulb = 1.12 / 22.4 =0.05
Final moles of H2 in 1st bulb = 0.05/2
No. of moles of H2 in 2nd bulb = 0.05/2 => Wt of H2 = 0.05 g

nD2 = 0.05/2*1.414 = 0.0176
=> Wt of D2 = 4 * 0.0176 = 0.0707 g

% of hydrogen=41.4%

(~ means proportional to)

Best Of luck

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert